A tank measuring 28 cm by 20 cm by 30 cm was filled with 5000 cm3 of water. At 2 p.m., water was continuously added into the tank at a constant rate of 25 cm3 every minute. At 3 p.m., 200 cm3 of water was drawn out of the tank and subsequently, a further 200 cm3 was also drawn out at the end of every hour after. At what time did the tank start to overflow for the first time? (Express your answer in HHMM format of the 24-hour clock. For example, 4 a.m. represented as 0400; 4.25 p.m. represented as 1625.)
thank you very much

Answer by ankor@dixie-net.com(22740)

A tank measuring 28 cm by 20 cm by 30 cm was filled with 5000 cm^3 of water.
Find how much water the tank holds when it's full.
28 * 30 * 30 = 16800 cu/cm when it is full
then find out how much more water required to fill the tank
16800 - 5000 = 11800
At 2 p.m., water was continuously added into the tank at a constant rate of 25 cm^3 every minute.
At 3 p.m., 200 cm^3 of water was drawn out of the tank and subsequently, a further 200 cm^3 was also drawn out at the end of every hour after.
60*25 = 1500 - 200 = 1300 cu/cm
At what time did the tank start to overflow for the first time?
1300 cu/cm is added every hour
find how many hours (h) it take to fill the tank at 1300 cu/cm per hr
1300h = 11800
h = 11800/1300
h = 9.08 hrs
But the last hr it won't get a chance to draw out the 200 cm^3
After 8 hrs: 5000 + 8(1300) = 15400 cm^3 in the tank,
16800 - 15400 = 1400 cu/3 space remains
Find no. of min (m) to fill this amt
25m = 1400
m = 1400/25
m = 56 min
Overflow should occur 8 hrs and 56 min from 2 PM
:
Find the time of overflow
1400 + 856 = 2256 is the time when overflow occurs